Calculate the momentum of a particle knowing that its wavelength is 10.13 x 10^-5 m.
1
Expert's answer
2020-05-15T09:02:08-0400

Given data

Wavelength is p=10.13x10-5 m

Planck's constant is h=6.626x10-34 J.s

----------------------------------------------------------------------------------------

According to de Broglie hypothesis, the wavelength of matter wave is

"\lambda=\frac{h}{p}"

The momentum of the particle is

"p=\frac{h}{\lambda}"

"=\frac{6.626*10^{-34} J.s}{10.13*10^{-5} m}"

"=6.54*10^{-30} kg.m\/s"